Jorge is taking a multiple choice test. He decided to guess on the first two problems. On each question there are 4 answer choices. What is the probability that he answers the first question correctly and the second question incorrectly?
Answer by stanbon(57224)

Jorge is taking a multiple choice test. He decided to guess on the first two problems. On each question there are 4 answer choices. What is the probability that he answers the first question correctly and the second question incorrectly?
----
The answers on the 2 questions are independent events.
P(correct followed by incorrect) = (1/4)(3/4) = 3/16
=========================
Cheers,
Stan H.